# Neuronal Attention Circuit (NAC) for Representation Learning

### v2

@@ Abstract
Attention improves representation learning over RNNs, but its discrete nature limits continuous-time (CT) modeling. We introduce Neuronal Attention Circuit (NAC), a novel, biologically plausible CT-Attention mechanism that reformulates attention logits computation as the solution to a linear first-order ODE with nonlinear interlinked gates derived from repurposing C. elegans Neuronal Circuit Policies (NCPs) wiring mechanism. NAC replaces dense projections with sparse sensory gates for key-query projections and a sparse backbone network with two heads for computing content-target and learnable time-constant gates, enabling efficient adaptive dynamics. NAC supports three attention logit computation modes: (i) explicit Euler integration, (ii) exact closed-form solution, and (iii) steady-state approximation. To improve memory intensity, we implemented a sparse Top-K pairwise concatenation scheme that selectively curates key-query interactions. We provide rigorous theoretical guarantees, including state stability, bounded approximation errors, and universal approximation. Empirically, we implemented NAC in diverse domains, including irregular time-series classification, lane-keeping for autonomous vehicles, and industrial prognostics. We observed that NAC matches or outperforms competing baselines in accuracy and occupies an intermediate position in runtime and memory efficiency compared with several CT baselines.
@@ 1 Introduction
Learning representations of sequential data in temporal or spatio-temporal domains is essential for capturing patterns and enabling accurate forecasting. Discrete-time Recurrent neural networks (DT-RNNs) such as RNN (rumelhart1985learning; jordan1997serial), Long-short term memory (LSTM) (hochreiter1997long), and Gated Recurrent Unit (GRU) (cho2014learning) model sequential dependencies by iteratively updating hidden states to represent or predict future elements in a sequence. While effective for regularly sampled sequences, DT-RNNs face challenges with irregularly sampled data because they assume uniform time intervals. In addition, vanishing gradients can make it difficult to capture long-term dependencies (hochreiter1998vanishing).
@@
Continuous-time RNNs (CT-RNNs) (rubanova2019latent) model hidden states as ordinary differential equations (ODEs), allowing them to process inputs that arrive at arbitrary or irregular time intervals. Mixed-memory RNNs (mmRNNs) (lechner2022mixed) build on this idea by separating memory compartments from time-continuous states, helping maintain stable error propagation while capturing continuous-time dynamics. Liquid neural networks (LNNs) (hasani2021liquid; hasani2022closed) take a biologically inspired approach by assigning variable time-constants to hidden states, improving adaptability and robustness, though vanishing gradients can still pose challenges during training.
@@
The attention mechanisms (vaswani2017attention) mitigate this limitation by treating all time steps equally and allowing models to focus on the most relevant observations. It computes the similarity between queries ($q$) and keys ($k$), scaling by the key dimension to keep gradients stable. Multi-Head Attention (MHA) (vaswani2017attention) extends this by allowing the model to attend to different representation subspaces in parallel. Variants like Sparse Attention (tay2020sparse; roy2021efficient), BigBird (zaheer2020big), and Longformer (beltagy2020longformer) modify the attention pattern to reduce computational cost, particularly for long sequences, by attending only to selected positions rather than all pairs. Even with these improvements, attention-based methods still rely on discrete scaled dot-product operations, limiting their ability to model continuous trajectories often captured by CT counterparts.
@@
Recent work has explored bridging this gap through Neural-ODE (chen2018neural) formulation. mTAN (shukla2021multi) learns CT embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. ODEFormer (d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to directly infer symbolic ODE systems from noisy, irregular data, though it struggles with chaotic systems and generalization beyond observed conditions. Continuous-time Attention (CTA) (chien2021continuous) embeds a continuous-time attention mechanism within a Neural ODE, allowing attention weights and hidden states to evolve jointly over time. Still, it remains computationally intensive and sensitive to the accuracy of the ODE solver. ContiFormer (chen2023contiformer) builds a CT-transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in data.
@@
Despite these innovations, a persistent and underexplored gap remains in developing a biologically plausible attention mechanism that seamlessly integrates CT dynamics with the abstraction of the brain’s connectome to handle irregular sequences without prohibitive computational costs. Building on this, we propose a novel attention mechanism called the Neuronal Attention Circuit (NAC), in which attention logits are computed as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing Neuronal Circuit Policies (NCPs) from the nervous system of C. elegans nematode (refer to Appendix A.2 for more information). Unlike standard attention, which projects key-query pairs through a dense layer, NAC employs a sensory gate to transform input features and a backbone to model nonlinear interactions, with multiple heads producing outputs structured for attention logits computation. Based on the solutions to ODE, we define three computation modes: (i) Exact, using the closed-form ODE solution; (ii) Euler, approximating the solution via explicit Euler integration; and (iii) Steady, using only the steady-state solution, analogous to standard attention scores. To reduce computational complexity, we implemented a sparse Top-K pairwise concatenation algorithm that selectively curates key-query inputs. We evaluate NAC across multiple domains, including irregularly sampled time series, autonomous vehicle lane-keeping, and Industry 4.0, comparing it to state-of-the-art baselines. NAC consistently matches or outperforms these models, while runtime and peak memory benchmarks place it between CT-RNNs in terms of speed and CT-Attentions in terms of memory requirements.
@@ 2 Neuronal Attention Circuit (NAC)
$\frac{da_{t}}{dt}=-\,\underbrace{f_{\omega_{\tau}}([\mathbf{q};\mathbf{k}],\theta_{\omega_{\tau}})}_{\omega_{\tau}(\mathbf{u})}a_{t}+\underbrace{f_{\phi}([\mathbf{q};\mathbf{k}],\theta_{\phi})}_{\phi(\mathbf{u})},$ (1)
where $\mathbf{u}=[\mathbf{q};\mathbf{k}]$ denotes the sparse Top-K concatenated query–key input. $\omega_{\tau}$ represents a learnable time-constant gate head, $\phi$ denotes a nonlinear content-target head. Both gates are parameterized by a backbone network derived from repurposing NCPs. We refer to this formulation as the Neuronal Attention Circuit (NAC). It enables the logits $a_{t}$ to evolve dynamically with input-dependent, variable time constants, mirroring the adaptive temporal dynamics found in C. elegans nervous systems while improving computational efficiency and expressiveness. Moreover, it introduces continuous depth into the attention mechanism, bridging discrete-layer computation with dynamic temporal evolution.
Motivation behind this formulation: The proposed formulation is loosely motivated by the input-dependent time-constant mechanism of Liquid Neural Networks (LNNs), a class of CT-RNNs inspired by biological nervous systems and synaptic transmission. In this framework, the dynamics of non-spiking neurons are described by a linear ODE with nonlinear, interlinked gates: $\frac{d\mathbf{x_{t}}}{dt}=\frac{\mathbf{x_{t}}}{\tau}+\mathbf{S_{t}},$ where $\mathbf{x_{t}}$ denotes the hidden state and $\mathbf{S_{t}}\in\mathbb{R}^{M}$ represents a nonlinear contribution defined as $f(\mathbf{x_{t}},\mathbf{u},t,\theta)(A-\mathbf{x_{t}})$. Here, $A$ and $\theta$ are learnable parameters. Plugging $\mathbf{S_{t}}$ yields $\frac{d\mathbf{x_{t}}}{dt}=-\left[-\frac{1}{\tau}-f(\mathbf{x_{t}},\mathbf{u},t,\theta)\right]\mathbf{x_{t}}+f(\mathbf{x_{t}},\mathbf{u},t,\theta)A$. LNNs are known for their strong expressivity, stability, and performance in irregularly sampled time-series modeling (hasani2021liquid; hasani2022closed).
@@
Closed-form (Exact) Computation of NAC: We now devise the analytical solution for Eqn. 1. Let both $\omega_{\tau}$ and $\phi$ be fixed in pseudo-time interval (frozen-coefficient approximation (john1952integration)) with initial condition $a_{0}$, then closed-form solution is:
@@
Here, $a^{*}=\phi/\omega_{\tau}$ is the steady-state solution. The full derivation is provided in Appendix B.1.
@@ 2.1.1 State Stability
We analyze state stability in both single-connection and multi-connection settings. This analysis establishes the boundedness of the attention logit state trajectory, ensuring that, under positive decay rates, the dynamics remain well-behaved without divergence or overshoot.
@@ 2.1.2 Closed-form Error & Exponential Bounds
Remark: If $\omega_{\tau}>0$ then $\lim_{t\to\infty}e^{-\omega_{\tau}t}=0$, therefore $a_{t}\to a^{*}$. The convergence is exponential with rate $\omega_{\tau}$. If $\omega_{\tau}<0$ then $e^{-\omega_{\tau}t}=e^{|\omega_{\tau}|t}$ diverges so $a_{t}$ grows exponentially away from $a^{*}$ in magnitude (unless initial offset $a_{0}-a^{*}=0$, a measure-zero case). If $\omega_{\tau}=0$ the ODE is $\dot{a}=\phi$ and the solution is linear in $t$ (unless $\phi=0$). For bounded dynamics that converge to an interpretable steady-state, it is required that $\omega_{\tau}>0$.
@@
If the initialization is only known to belong to a bounded set, i.e., $|a_{0}-a^{*}|\leq M$ for some $M>0$, then the error admits the uniform bound
@@
Remark: The convergence rate is inversely proportional to $\omega_{\tau}$, and the required time scales only logarithmically in the accuracy level $1/\delta$. Intuitively, larger $\omega_{\tau}$ accelerates contraction towards equilibrium, yielding faster attainment of any prescribed tolerance.
Figure 1: Illustration of (a) NCPs with pre-determined wiring; (b) Sensory gate, where sensory neurons are active, and the remaining neurons are disabled for the $q$, $k$, and $v$ projections; (c) Backbone, showing inter-motor projections with sensory neurons disabled in extended heads for computing $\phi$ and $\omega_{\tau}$.
@@
0:  Keys $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$, Top-K value $K$
@@
  Scores: $S\leftarrow Q\cdot K^{\top}$
  Effective Top-K: $K_{\text{eff}}\leftarrow\min(K,T_{k})$
  Indices: $I_{\text{topk}}\leftarrow\text{top\_k}(S,K_{\text{eff}})$
  Gather: $K_{\text{selected}}\leftarrow\text{gather}(K,I_{\text{topk}})\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times D}$
  Tiled: $Q_{\text{tiled}}\leftarrow\text{tile}(Q,K_{\text{eff}})\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times D}$
  Concatenate: $U_{\text{topk}}\leftarrow[\,Q_{\text{tiled}};\;K_{\text{selected}}\,]\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$
@@ 2.2 Designing the Neural Network
Repurposing NCPs: We repurpose the NCPs framework by converting its fixed, biologically derived wiring (see Figure 1(a)) into a flexible recurrent architecture that allows configurable input–output mappings. Instead of enforcing a static connectome, our approach exposes adjacency matrices as modifiable structures defining sparse input and recurrent connections. This enables selective information routing across neuron groups while retaining the original circuit topology. Decoupling wiring specifications from model instantiation allows dynamic connectivity adjustments to accommodate different input modalities without full retraining. Algorithm 1 summarizes the steps for repurposing the NCPs wiring mechanism. Key features include group-wise masking for neuron isolation, adaptive remapping of inputs and outputs for task-specific adaptation, and tunable sparsity $s$ to balance expressiveness and efficiency.
In our implementation, the sensory neuron gate ($\mathcal{NN}_{\text{sensory}}$) projects the $q$, $k$, and $v$ representations (see Figure 1(b)). This enables sensory neurons to maintain structured, context-aware representations rather than collapsing inputs into fully connected layers. As a result, the network preserves locality and modularity, which improves information routing.
$\mathcal{NN}_{\text{sensory}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{s}],\,G_{\text{output}}=[N_{s}],$ (13)
$\qquad\quad\mathcal{D}=[N_{i},N_{c},N_{m}],\,s)$
The inter-to-motor pathways form a backbone network ($\mathcal{NN}_{\text{backbone}}$) with branches that compute $\phi$ and $\omega_{\tau}$ (see Figure 1(c)). Instead of learning $\phi$ and $\omega_{\tau}$ independently, this backbone allows the model to learn shared representations, enabling multiple benefits: (i) separate head layers enable the system to capture temporal and structural dependencies independently; (ii) accelerates convergence during training.
$\mathcal{NN}_{\text{backbone}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{i}],\,G_{\text{output}}=[N_{m}],$ (14)
$\qquad\quad\mathcal{D}=[N_{s}],\,s)$
@@
Input Curation: We experimented with different strategies for constructing query–key inputs. Initially, we implemented full pairwise concatenation, where queries $Q\in\mathbb{R}^{B\times H\times T_{q}\times D}$ are combined with all keys $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$ to form a joint tensor $U\in\mathbb{R}^{B\times H\times T_{q}\times T_{k}\times 2D}$. While this preserved complete feature information and enabled expressive, learnable similarity functions, it was memory-intensive, making it impractical for longer sequences. To mitigate this, we applied a sparse Top-K optimization: for each query, we compute pairwise scores $S=Q\cdot K^{\top}\in\mathbb{R}^{B\times H\times T_{q}\times T_{k}}$, select the Top-$K_{\text{eff}}=\min(K,T_{k})$ keys, and construct concatenated pairs $U_{\text{topk}}\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$. This approach preserves the most relevant interactions while substantially reducing memory requirements in the concatenation and subsequent backbone processing stages, allowing the method to scale linearly with the sequence length in those components. However, the initial computation of $S$ remains quadratic (see Appendix C.3). Algorithm 2 outlines the steps required for input curation.
Time Vector: NAC builds on continuous-depth models as (hasani2022closed) that adapt their temporal dynamics to the task. It constructs an internal, normalized pseudo-time vector $t_{\text{pseudo}}$ using a sigmoidal transformation, $t_{\text{pseudo}}=\sigma(t_{a}\cdot t+t_{b})$, where $t_{a}$ and $t_{b}$ are learnable affine parameters and $\sigma$ is the sigmoid function. For time-varying datasets (e.g., irregularly sampled series), each time point $t$ is derived from the sample’s timestamp, while for tasks without meaningful timing, $t$ is set to $1$. The resulting $t_{\text{pseudo}}$ lies in $[0,1]$ and provides a smooth, bounded representation of time for modulating the network’s dynamics.
@@
For finite $t$, the exponential factor $(1-e^{-\omega_{\tau}t})$ regulates the buildup of attention, giving $\omega_{\tau}$ a temporal gating role. Normalizing across all keys via softmax yields attention weights $\alpha_{t}=\text{softmax}(a_{t})$, defining a valid probability distribution where $\phi$ amplifies or suppresses content alignments, and $\omega_{\tau}$ shapes both the speed and saturation of these preferences.
As $t\to\infty$, the trajectory converges to the steady state
$a^{*}_{t}=\frac{\phi}{\omega_{\tau}}\approx\frac{q^{\top}k}{\sqrt{d_{k}}},$ (18)
which is analogous to scaled-dot attention under specific parameterization when the backbone $\mathcal{NN}_{\text{backbone}}$ is configured as a linear projection such that $\phi(\mathbf{u})=q^{\top}k$ and $\omega_{\tau}(u)=\sqrt{d_{k}}$ (e.g., by setting NCP weights to emulate bilinear forms and disabling nonlinearities). In general, the nonlinear backbone allows for more expressive similarities, with the approximation holding when trained to mimic dot products.
Attention output: Finally, the attention output is computed by integrating the attention weights with the value matrix:
@@
In practice, the integration is approximated using a Riemann-style approach, where the weighted elements are computed by multiplying each $v_{t}$ with its corresponding $\alpha_{t}$. These are then summed and multiplied by a fixed pseudo-time step $\delta_{t}$, chosen as a scalar (typically between 0.5–1.0) hyperparameter during layer initialization. This yields a continuous analogue of standard weighted sums, giving finer resolution of the attention trajectory without altering the underlying values. Sensitivity to attention output w.r.t $\delta_{t}$ is visualized in Appendix D.2.
2.2.1 Extension to Multi-head
To scale this mechanism to multi-head attention, we project the input sequence into $H$ independent subspaces (heads) of dimension $d_{\text{model}}/H$, yielding query, key, and value tensors $(q^{(h)},k^{(h)},v^{(h)})$ for $h\in\{1,\dots,H\}$. For each head, pairwise logits are computed according to Eqns. 2,3 or 18, followed by the softmax normalization to calculate attention weights. The resulting attention weights $\alpha^{(h)}_{t}$ are then used to integrate with the value vector $v^{(h)}$, producing head-specific attention outputs. Finally, these outputs are concatenated and linearly projected back into the model dimension. This formulation ensures that each head learns distinct dynamic compatibilities governed by its own parameterization of $\phi$ and $\omega_{\tau}$, while the aggregation across heads preserves the expressive capacity of the standard multi-head attention mechanism.

### v3

@@ Abstract
Attention improves representation learning over RNNs, but its discrete nature limits continuous-time (CT) modeling. We introduce Neuronal Attention Circuit (NAC), a novel, biologically inspired CT-Attention mechanism that reformulates attention logit computation as the solution to a linear first-order ODE with nonlinear interlinked gates derived from repurposing C.elegans Neuronal Circuit Policies (NCPs) wiring. NAC replaces dense projections with sparse sensory gates for key-query projections and a sparse backbone network with two heads for computing content-target and learnable time-constant gates, enabling efficient adaptive dynamics. To improve efficiency and memory consumption, we implemented an adaptable subquadratic sparse Top-K pairwise concatenation mechanism that selectively curates key-query interactions. We provide rigorous theoretical guarantees, including state stability and bounded approximation errors. Empirically, we implemented NAC in diverse domains, including irregular time-series classification, lane-keeping for autonomous vehicles, and industrial prognostics. We observed that NAC matches or outperforms competing baselines in accuracy and occupies an intermediate position in runtime and memory consumption compared with several CT state-of-the-art baselines, while being interpretable at the neuron cell level.
@@ 1 Introduction
Learning representations of sequential data in temporal or spatiotemporal domains is essential for capturing patterns and enabling accurate forecasting. Discrete-time Recurrent neural networks (DT-RNNs), such as RNN (rumelhart1985learning; jordan1997serial), Long-short term memory (LSTM) (hochreiter1997long), and Gated Recurrent Unit (GRU) (cho2014learning) model sequential dependencies by iteratively updating hidden states to represent or predict future elements in a sequence. While effective for regularly sampled sequences, DT-RNNs face challenges with irregularly sampled data because they assume uniform time intervals. In addition, vanishing gradients can make it difficult to capture long-term dependencies (hochreiter1998vanishing).
@@
Continuous-time RNNs (CT-RNNs) (rubanova2019latent) model hidden states as ordinary differential equations (ODEs), allowing them to process inputs that arrive at arbitrary or irregular time intervals. Mixed-memory RNNs (mmRNNs) (lechner2022mixed) build on this idea by separating memory compartments from time-continuous states, helping maintain stable error propagation while capturing continuous-time dynamics. Liquid neural networks (LNNs) (hasani2021liquid; hasani2022closed) take a biologically inspired approach by assigning variable time-constants to hidden states, improving adaptability and robustness, although vanishing gradients can still pose challenges during training.
@@
The attention mechanisms (vaswani2017attention) mitigate this limitation by treating all time steps equally and allowing models to focus on the most relevant observations. It computes the similarity between queries ($q$) and keys ($k$), scaling by the key dimension to keep the gradients stable. Multihead Attention (MHA) (vaswani2017attention) extends this by allowing the model to attend to different representation subspaces in parallel. Variants such as Sparse Attention (zhang2021sparse), BigBird (zaheer2020big), and Longformer (beltagy2020longformer) modify the attention pattern to reduce computational cost, particularly for long sequences, by attending only to selected positions rather than all pairs. Even with these improvements, the computation paths in these methods are discrete in nature, limiting their ability to model continuous trajectories often captured by their CT counterparts.
@@
Recent work has explored bridging this gap through the Neural-ODE (chen2018neural) formulation. mTAN (shukla2021multi) learns CT embeddings and uses time-based attention to interpolate irregular observations into a fixed-length representation for downstream encoder-decoder modeling. Continuous-time Attention (CTA) (chien2021continuous) embeds a CT-Attention mechanism within a Neural ODE, allowing attention weights and hidden states to evolve jointly over time. Nevertheless, it remains computationally intensive and sensitive to the accuracy of the ODE solver. ODEFormer (d2023odeformer) trains a sequence-to-sequence transformer on synthetic trajectories to infer symbolic ODE systems directly from noisy, irregular data, although it struggles with chaotic systems and generalization beyond observed conditions. ContiFormer (chen2023contiformer) builds a CT-Transformer by pairing ODE-defined latent trajectories with a time-aware attention mechanism to model dynamic relationships in data.
@@
Despite these advancements, a persistent and underexplored gap remains in developing a biologically inspired attention mechanism that seamlessly integrates CT dynamics with the abstraction of the brain’s connectome. Building on this, we propose a novel attention mechanism called the Neuronal Attention Circuit (NAC), in which attention logits are computed as the solution to a first-order ODE modulated by nonlinear, interlinked gates derived from repurposing Neuronal Circuit Policies (NCPs) from the nervous system of C.elegans nematode (refer to Appendix A.1 for more information). Unlike standard attention, which projects key-query pairs through a dense layer, NAC employs a sensory gate to transform input features and a backbone network to model nonlinear interactions, with multiple heads producing outputs structured for attention logit computation. Based on the solutions to ODE, we define three computation modes: (i) Exact, using the closed-form ODE solution; (ii) Euler, approximating the solution via explicit Euler integration; and (iii) Steady, using only the steady-state solution, analogous to standard attention scores. To reduce computational complexity, we implement a sparse subquadratic Top-K pairwise concatenation algorithm that selectively curates key-query inputs. We evaluate NAC across multiple domains, including irregularly sampled time-series, autonomous vehicle lane-keeping, and Industry 4.0, and compare it to state-of-the-art baselines. NAC consistently matches or outperforms these models, while runtime and peak memory benchmarks place it between CT-RNNs in terms of speed and CT-Attentions in terms of memory requirements. We also highlight the interpretability of NAC at the cell level.
@@ 2 Neuronal Attention Circuit (NAC)
$\frac{da_{t}}{dt}=-\,\underbrace{f_{\omega_{\tau}}([\mathbf{q};\mathbf{k}],t,\theta_{\omega_{\tau}})}_{\omega_{\tau}(\mathbf{u})}a_{t}+\underbrace{f_{\phi}([\mathbf{q};\mathbf{k}],t,\theta_{\phi})}_{\phi(\mathbf{u})},$ (1)
where $\mathbf{u}=[\mathbf{q};\mathbf{k}]$ denotes the sparse Top-K concatenated query–key input. $\omega_{\tau}$ represents a learnable time-constant gate head, and $\phi$ denotes a nonlinear content-target head. Both gates are parameterized by a backbone network derived from repurposing NCPs. We refer to this formulation as the Neuronal Attention Circuit (NAC). It enables the logit $a_{t}$ to evolve dynamically with input dependent, variable time constants, mirroring the adaptive temporal dynamics found in C.elegans nervous systems while improving computational efficiency and expressiveness. Moreover, it introduces continuous depth into the attention mechanism, bridging discrete-layer computation with dynamic temporal evolution.
Motivation behind this formulation: The proposed formulation is loosely motivated by the input-dependent time-constant mechanism of LNNs, a class of CT-RNNs inspired by biological nervous systems and synaptic transmission. In this framework, the dynamics of nonspiking neurons are described by a linear ODE with nonlinear, interlinked gates: $\frac{d\mathbf{x_{t}}}{dt}=\frac{\mathbf{x_{t}}}{\tau}+\mathbf{S_{t}},$ where $\mathbf{x_{t}}$ denotes the hidden state and where $\mathbf{S_{t}}\in\mathbb{R}^{M}$ represents a nonlinear contribution defined as $f(\mathbf{x_{t}},\mathbf{u},t,\theta)(A-\mathbf{x_{t}})$. Here, $A$ and $\theta$ are learnable parameters. Plugging $\mathbf{S_{t}}$ yields $\frac{d\mathbf{x_{t}}}{dt}=-\left[\frac{1}{\tau}+f(\mathbf{x_{t}},\mathbf{u},t,\theta)\right]\mathbf{x_{t}}+f(\mathbf{x_{t}},\mathbf{u},t,\theta)A$. LNNs are known for their strong expressivity, stability, and performance in irregularly sampled time-series modeling (hasani2021liquid; hasani2022closed).
@@
Closed-form (Exact) Computation of NAC: We now derive the analytical solution for Eqn. 1. The gates $\phi$ and $\omega_{\tau}$ are time-varying and depend on the input sequence at each time step. However, to enable an efficient closed-form forward-pass update, we apply a piecewise constant assumption (locally frozen-coefficient approximation (john1952integration)) over a small interval. Under this assumption, with initial condition $a_{0}$, the analytical closed-form solution is as follows:
@@
Here, $a^{*}=\phi/\omega_{\tau}$ is the steady-state solution. The detailed derivation and analysis are provided in Appendix B.1.
@@ 2.1.1 State Stability
We analyze state stability in both single-connection and multi-connection settings. This analysis establishes the boundness of the attention logit state trajectory, ensuring that, under positive decay rates, the dynamics remain well-behaved without divergence or overshoot.
@@ 2.1.2 Closed-form Error & Exponential Bounds
Remark: If $\omega_{\tau}>0$ then $\lim_{t\to\infty}e^{-\omega_{\tau}t}=0$, therefore $a_{t}\to a^{*}$. The convergence is exponential with rate $\omega_{\tau}$. If $\omega_{\tau}<0$ then $e^{-\omega_{\tau}t}=e^{|\omega_{\tau}|t}$ diverges so that $a_{t}$ grows exponentially away from $a^{*}$ in magnitude (unless initial offset $a_{0}-a^{*}=0$, a measure-zero case). If $\omega_{\tau}=0$ the ODE is $\dot{a}=\phi$ and the solution is linear in $t$ (unless $\phi=0$). For bounded dynamics that converge to an interpretable steady-state, it is required that $\omega_{\tau}>0$.
@@
If the initialization is known only to belong to a bounded set, i.e., $|a_{0}-a^{*}|\leq M$ for some $M>0$, then the error admits the uniform bound
@@
Remark: The convergence rate is inversely proportional to $\omega_{\tau}$, and the required time scales only logarithmically at the accuracy level $1/\delta$. Intuitively, a larger $\omega_{\tau}$ accelerates contraction toward equilibrium, yielding faster attainment of any prescribed tolerance.
Figure 1: Illustration of (a) NCPs with predetermined wiring; (b) Sensory gate, where sensory neurons are active, and the remaining neurons are disabled for the $q$, $k$, and $v$ projections; (c) Backbone, showing inter-motor projections with sensory neurons disabled in extended heads for computing $\phi$ and $\omega_{\tau}$.
@@
0:  Queries $Q\in\mathbb{R}^{B\times H\times T_{q}\times D}$, Keys $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$, Top-K value $K$
@@
  Block Size: $B_{s}\leftarrow\lfloor\sqrt{T_{k}}\rfloor$
  Partition $K$ into $N_{b}$ blocks: $K_{\text{blocks}}\in\mathbb{R}^{B\times H\times N_{b}\times B_{s}\times D}$
  Block Centroids: $\bar{K}\leftarrow\text{mean}(K_{\text{blocks}})$
  Coarse Scores: $S_{\text{coarse}}\leftarrow Q\cdot\bar{K}^{\top}$
  Select Top-$M$ Blocks: $I_{\text{blocks}}\leftarrow\text{top\_k}(S_{\text{coarse}},\lceil K/B_{s}\rceil)$
  Gather Candidates: $K_{\text{cand}}\leftarrow\text{gather}(K_{\text{blocks}},I_{\text{blocks}})$
  Fine Scores: $S_{\text{fine}}\leftarrow\text{reduce\_sum}(Q\odot K_{\text{cand}})$
  Effective Top-K: $K_{\text{eff}}\leftarrow\min(K,\text{size}(K_{\text{cand}}))$
  Indices: $I_{\text{topk}}\leftarrow\text{top\_k}(S_{\text{fine}},K_{\text{eff}})$
  Selected Keys: $K_{\text{selected}}\leftarrow\text{gather}(K_{\text{cand}},I_{\text{topk}})$
  Concatenate: $U_{\text{topk}}\leftarrow[\,Q;\;K_{\text{selected}}\,]\in\mathbb{R}^{B\times H\times T_{q}\times K_{\text{eff}}\times 2D}$
@@ 2.2 Designing the Neural Network
Repurposing NCPs: We repurpose the NCPs framework by converting its fixed, biologically derived wiring (see Figure 1(a)) into a flexible recurrent architecture that allows configurable input–output mappings. Instead of enforcing a static connectome, our approach exposes adjacency matrices as modifiable structures defining sparse input and recurrent connections. This enables selective information routing across neuron groups while retaining the original circuit topology. Decoupling wiring specifications from model instantiation allows dynamic connectivity adjustments to accommodate different input modalities without full retraining. Algorithm 1 summarizes the steps for repurposing the NCPs wiring mechanism. Key features include groupwise masking for neuron isolation, adaptive remapping of inputs and outputs for task-specific adaptation, and tunable sparsity $s$ to balance expressiveness and efficiency.
In our implementation, the sensory neuron gate ($\mathcal{NN}_{\text{sensory}}$) projects the $q$, $k$, and $v$ representations (see Figure 1(b)). This enables sensory neurons to maintain structured, context-aware representations rather than collapsing inputs into fully connected layers. As a result, the network preserves locality and modularity, which improves information routing. Intuitively, it acts like biological receptor neurons, converting external stimulus inputs into organized signals that feed into downstream circuits.
$\mathcal{NN}_{\text{sensory}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{s}],\,$ (13)
$\qquad G_{\text{output}}=[N_{s}],\mathcal{D}=[N_{i},N_{c},N_{m}],\,s)$
The inter-to-motor pathways form a backbone network ($\mathcal{NN}_{\text{backbone}}$) with branches that compute $\phi$ and $\omega_{\tau}$ (see Figure 1(c)). Intuitively, this integrates neuronal activity over time and produces outputs that regulate gating signals. Instead of learning $\phi$ and $\omega_{\tau}$ independently, this backbone allows the model to learn shared representations, enabling multiple benefits: (i) separate head layers enable the system to capture temporal and structural dependencies independently; (ii) accelerates convergence during training.
$\mathcal{NN}_{\text{backbone}}$ $=\text{NCPCell}(G_{\text{input}}=[N_{i}],\$(14)
$\qquad G_{\text{output}}=[N_{m}],\mathcal{D}=[N_{s}],\,s)$
@@
Input Curation: Initially, full pairwise concatenation was used to form a joint tensor $U\in\mathbb{R}^{B\times H\times T_{q}\times T_{k}\times 2D}$ from $Q\in\mathbb{R}^{B\times H\times T_{q}\times D}$ and $K\in\mathbb{R}^{B\times H\times T_{k}\times D}$. However, this approach becomes computationally intractable as the sequence length increases, leading to $\mathcal{O}(n^{2})$ memory growth, severe memory fragmentation, and frequent hardware memory exhaustion. A naive sparse Top-K pruning strategy was then implemented to reduce the load on subsequent layers. While this reduced downstream computation, it failed to address the fundamental quadratic bottleneck in the initial scoring and storage stages, rendering the optimization ineffective for long sequences. To mitigate this issue, we then implemented a subquadratic Top-K selection strategy based on square-root partitioning  (child2019sparse). By dynamically segmenting the key sequence into blocks and scoring queries against block centroids, the method identifies high-salience candidate regions without ever materializing the full $T_{q}\times T_{k}$ matrix. This hierarchical approach reduces the scoring complexity to $\mathcal{O}(n\sqrt{n})$, producing a sparse tensor $U_{\text{topk}}$ that enables linear memory scaling in subsequent backbone layers. The mechanism is self-adaptive and eliminates the need for manual block-size tuning (Algorithm 2). The efficiency and complexity comparison is provided in Appendix C.2.
Construction of Time Vector ($t$): NAC builds on continuous-depth models as in (hasani2022closed), which adapt their temporal dynamics to the task. It constructs an internal, normalized time vector $t=\sigma(t_{a}\cdot t_{\text{sample}}+t_{b})\in[0,T]$, where $t_{a}$ and $t_{b}$ are learnable affine parameters (computed per query-key pair) and $\sigma$ is the sigmoid function. For time-varying datasets (e.g., irregularly sampled series), each $t_{\text{sample}}$ is derived from the sample’s timestamp, enabling time-dependent modulation of attention dynamics. For tasks without meaningful timing, $t_{\text{sample}}$ is set to $1$, resulting in $t=\sigma(t_{a}+t_{b})\in[0,1]$, providing a smooth, bounded representation of $t$ for modulating the network’s dynamics.
@@
For finite $t_{\text{sample}}$, the exponential factor $(1-e^{-\omega_{\tau}t})$ regulates the buildup of attention, giving $\omega_{\tau}$ a temporal gating role. Normalizing across all keys via softmax yields attention weights $\alpha_{t}=\text{softmax}(a_{t})$, defining a valid probability distribution where $\phi$ amplifies or suppresses content alignments, and $\omega_{\tau}$ shapes both the speed and saturation of these preferences.
As $t_{\text{sample}}\to\infty$, the trajectory converges to the steady state
$a^{*}_{t}=\frac{\phi}{\omega_{\tau}},$ (18)
which we introduce as a computation mode. Conceptually, it is analogous to scaled-dot attention: it aggregates information across inputs in a similar form, even though $\phi$ cannot reproduce exact dot products.
Attention output: Finally, the attention output is computed by integrating the $\alpha_{t}$ with the $v_{t}$ matrix over internally constructed time-window $t$:
@@
Intuitively, this integral acts as a continuous aggregator of the $v_{t}$, where each contribution is weighted by its instantaneous relevance. In practice, we approximate it via a time-modulated sum: $\sum_{i}\alpha^{(i)}_{t}v^{(i)}_{t}\,t^{(i)}$ where $t^{(i)}$ serves as a flexible quadrature weight that adaptively scales the contribution of each time-step, complementing $\alpha^{(i)}_{t}$. This parameterization allows the model to modulate temporal contributions independently of semantic alignment while preserving differentiability and maintaining the original value representations.